# Neural Empirical Interpolation Method for Nonlinear Model Reduction

### v1

@@ 4.4. Q-Tensor Model of Liquid Crystals Finite Elements
where the unknown are the $2\times 2$ symmetric and trace-free Q-tensor and the scalar auxiliary variable $r$, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product, and
@@
Because $Q$ is symmetric and trace-free [17], this problem only depends on three variables: $Q_{11}$, $Q_{12}$, and $r$. To recast the problem in the ROM setting, we consider a parameterized nonlinear term where $a\in[-0.5,0.5]$ is the varying parameter while holding $M,L,b,c,A_{0}$ fixed. We set these values to be
$M=1,\quad L=0.1,\quad b=0.5,\quad c=1,\quad A_{0}=500,$
and we use homomgeneous Dirichlet boundary conditions and initial condition
$Q_{0}(x,y)=dd^{\top}-\frac{\|d\|_{2}^{2}}{2}I,\quad d=\frac{\begin{bmatrix}(2-x)(2+x)(2-y)(2+y)&\sin(\pi x)\sin(\pi y)\end{bmatrix}^{\top}}{\sqrt{1+[(2-x)(2+x)(2-y)(2+y)]^{2}+\sin(\pi x)^{2}\sin(\pi y)^{2}}},$
@@
$Q_{t}=M\left(L\Delta Q-\frac{\delta\mathcal{F}_{B}(Q)}{\delta Q}\right)=M\left(L\Delta Q-\left(aQ-b\left(Q^{2}-\frac{1}{2}\operatorname{tr}(Q^{2})I\right)+c\operatorname{tr}(Q^{2})Q\right)\right),$
so we can think of the parameter $a$ as the coefficient of the quadratic term in the quartic double-well bulk potential
$\mathcal{F}_{B}(Q)=\frac{a}{2}\operatorname{tr}(Q^{2})-\frac{b}{3}\operatorname{tr}(Q^{3})+\frac{c}{4}(\operatorname{tr}(Q^{2}))^{2}.$
Changing the parameter $a$ affects the depth of the two wells in the double-well potential while keeping the end behavior of the potential the same.
Due to the finite element discretization we exploit, and the complexity of the model, we need to approximate two different nonlinear terms via DEIM/NEIM to recover the solution of the problem. Denote by $\boldsymbol{Q}_{11}^{n}$, $\boldsymbol{Q}_{12}^{n}$, and $\boldsymbol{r}^{n}$ the vectors of solutions at time $n\Delta t$, with $\Delta t$ the time step.
@@
where $\Delta x_{i}$ is the volume of the finite element basis test function at node $i$. The second approximation is for
$f_{2}(\boldsymbol{Q}_{11}^{n},\boldsymbol{Q}_{12}^{n},\boldsymbol{Q}_{11}^{n+1},\boldsymbol{Q}_{12}^{n+1})_{i}=\left\langle P\left(\begin{bmatrix}(\boldsymbol{Q}_{11}^{n})_{i}&(\boldsymbol{Q}_{12}^{n})_{i}\\ (\boldsymbol{Q}_{12}^{n})_{i}&-(\boldsymbol{Q}_{11}^{n})_{i}\end{bmatrix}\right),\begin{bmatrix}(\boldsymbol{Q}_{11}^{n+1})_{i}-(\boldsymbol{Q}_{11}^{n})_{i}&(\boldsymbol{Q}_{12}^{n+1})_{i}-(\boldsymbol{Q}_{12}^{n})_{i}\\ (\boldsymbol{Q}_{12}^{n+1})_{i}-(\boldsymbol{Q}_{12}^{n})_{i}&-(\boldsymbol{Q}_{11}^{n+1})_{i}+(\boldsymbol{Q}_{11}^{n})_{i}\end{bmatrix}\right\rangle_{F}$
for $i=1,\dots,n$. The weights in NEIM are chosen as $w_{e}(\mu_{i};\mu_{j})=\delta_{ij}$ and
$w_{t}^{(k)}(a_{i},t_{i})=\begin{cases}1&\text{if $a_{i}=a_{k}$ and $|i-k|\leq 5$},\\ 0&\text{otherwise}.\end{cases}$
@@
We plot the behavior of the solutions in Figure 10. The blue vectors in the figure denote the orientation of the liquid crystal molecules at a given point, also called the director. In particular, a solution $Q(t,x,y)$ represents a distribution of orientations of liquid crystal molecules at the point $(x,y)$, so there is an average of that distribution and a measure of how aligned molecules in that distribution are with the average. If $\lambda_{1}(t,x,y)\geq\lambda_{2}(t,x,y)=-\lambda_{1}(t,x,y)$ are the eigenvalues of $Q(t,x,y)$ and $v_{1}(t,x,y),v_{2}(t,x,y)$ are the corresponding eigenvectors, then the blue vectors are the eigenvector $v_{1}(t,x,y)$. Then the orange contour in the figure denotes how much light passes through the liquid crystal at a given point in space, which is determined by the average orientation of the liquid crystal molecules $v_{1}(t,x,y)$ and the measure of alignment of molecules in the distribution with the average, $\lambda_{1}(t,x,y)$. Specifically, the color is given by a color mapping with the scalar quantity
$\lvert\lambda_{1}(t,x,y)(v_{1}(t,x,y)\cdot(1,0))\rvert,$
so if the molecules are in the vertical direction or are in the isotropic liquid phase (that is, they aren’t aligned with the average orientation so that $\lambda_{1}(t,x,y)\approx 0$) then light doesn’t pass through the liquid crystal. We see that qualitatively, the high fidelity, DEIM, and NEIM solutions all have similar contour plots, and the blue director fields are also similar. The contour for DEIM appears more similar to the high fidelity contour than NEIM compared to the high fidelity contour for time $t=0.375$, but NEIM is more similar at the final time $t=2$. In both cases, the blue director fields in the DEIM and NEIM plots well-approximate the high fidelity director field, with each plot in the equilibrium state of all horizontal directors.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. MH also acknowledges the ThinkSwiss Research Scholarship and the EPFL Excellence Research Scholarship. FP acknowledges the “GO for IT” program within the CRUI fund for the project “Reduced order method for nonlinear PDEs enhanced by machine learning”. This work has been conducted within the research activities of the consortium iNEST (Interconnected North-East Innovation Ecosystem), Piano Nazionale di Ripresa e Resilienza (PNRR) – Missione 4 Componente 2, Investimento 1.5 – D.D. 1058 23/06/2022, ECS00000043, supported by the European Union’s NextGenerationEU program.

### v2

@@ 4.4. Q-Tensor Model of Liquid Crystals Finite Elements
where the unknowns are the $2\times 2$ symmetric and trace-free Q-tensor and the scalar auxiliary variable $r$, $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product, and
@@
Because $Q$ is symmetric and trace-free [19], this problem only depends on three variables: $Q_{11}$, $Q_{12}$, and $r$. To recast the problem in the ROM setting, we consider a parameterized nonlinear term where $a\in[-0.5,0.5]$ is the varying parameter while holding the other parameters fixed at $M=1$, $L=0.1$, $b=0.5$, $c=1$, $A_{0}=500$, and we use homogeneous Dirichlet boundary conditions, and initial condition
$Q_{0}(x,y)=dd^{\top}-\frac{\|d\|_{2}^{2}}{2}I,\quad d=\frac{\begin{bmatrix}(4-x^{2})(4-y^{2})&\sin(\pi x)\sin(\pi y)\end{bmatrix}^{\top}}{\sqrt{1+[(4-x^{2})(4-y^{2})]^{2}+\sin(\pi x)^{2}\sin(\pi y)^{2}}},$
@@
$Q_{t}=M\left(L\Delta Q-\frac{\delta\mathcal{F}_{B}(Q)}{\delta Q}\right),$
with $\frac{\delta\mathcal{F}_{B}(Q)}{\delta Q}=aQ-b\left(Q^{2}-\frac{1}{2}\operatorname{tr}(Q^{2})I\right)+c\operatorname{tr}(Q^{2})Q$, so we can think of the parameter $a$ as the coefficient of the quadratic term in the quartic double-well bulk potential $\mathcal{F}_{B}(Q)=\frac{a}{2}\operatorname{tr}(Q^{2})-\frac{b}{3}\operatorname{tr}(Q^{3})+\frac{c}{4}(\operatorname{tr}(Q^{2}))^{2}$. Changing the parameter $a$ affects the depth of the two wells in the double-well potential while keeping the end behavior of the potential the same.
Due to the finite element discretization we exploit, and the complexity of the model, we need to approximate two different nonlinear terms via DEIM/NEIM to recover the solution of the problem. Indeed, the time discretization for the problem is given by
$\begin{cases}\frac{Q^{n+1}-Q^{n}}{\Delta t}&=M\left(\frac{L}{2}(\Delta Q^{n+1}+\Delta Q^{n})-r^{n}P(Q^{n})\right)\\ r^{n+1}-r^{n}&=\langle P(Q^{n}),Q^{n+1}-Q^{n}\rangle_{F},\end{cases}$
so we see that both $r^{n}P(Q^{n})$ and $\langle P(Q^{n}),Q^{n+1}-Q^{n}\rangle_{F}$ must be approximated. Denote by $\boldsymbol{Q}_{11}^{n}$, $\boldsymbol{Q}_{12}^{n}$, and $\boldsymbol{r}^{n}$ the vectors of solutions at time $n\Delta t$, with $\Delta t$ the time step.
@@
where $\Delta x_{i}$ is the volume of the finite element basis test function at node $i$.
The second approximation is for
$f_{2}(\boldsymbol{Q}_{11}^{n},\boldsymbol{Q}_{12}^{n},\boldsymbol{Q}_{11}^{n+1},\boldsymbol{Q}_{12}^{n+1})_{i}=$
$\left\langle P\left(\begin{bmatrix}(\boldsymbol{Q}_{11}^{n})_{i}&(\boldsymbol{Q}_{12}^{n})_{i}\\ (\boldsymbol{Q}_{12}^{n})_{i}&-(\boldsymbol{Q}_{11}^{n})_{i}\end{bmatrix}\right),\begin{bmatrix}(\boldsymbol{Q}_{11}^{n+1})_{i}-(\boldsymbol{Q}_{11}^{n})_{i}&(\boldsymbol{Q}_{12}^{n+1})_{i}-(\boldsymbol{Q}_{12}^{n})_{i}\\ (\boldsymbol{Q}_{12}^{n+1})_{i}-(\boldsymbol{Q}_{12}^{n})_{i}&-(\boldsymbol{Q}_{11}^{n+1})_{i}+(\boldsymbol{Q}_{11}^{n})_{i}\end{bmatrix}\right\rangle_{F}$
for $i=1,\dots,n$. The weights in NEIM are chosen as $w_{e}(\mu_{i};\mu_{j})=\delta_{ij}$ and, denoting $\mu_{i}=(a_{i},t_{i})$,
$w_{t}^{(k)}(a_{i},t_{i})=\begin{cases}1&\text{if $a_{i}=a_{k}$ and {$|t_{i}-t_{k}|\leq 5\Delta t$}},\\ 0&\text{otherwise},\end{cases}$
@@
We plot the behavior of the solutions in Figure 13. The blue vectors in the figure denote the orientation of the liquid crystal molecules at a given point, also called the director. In particular, a solution $Q(t,x,y)$ represents a distribution of orientations of liquid crystal molecules at the point $(x,y)$, so there is an average of that distribution and a measure of how aligned molecules in that distribution are with the average. If $\lambda_{1}(t,x,y)\geq\lambda_{2}(t,x,y)=-\lambda_{1}(t,x,y)$ are the eigenvalues of $Q(t,x,y)$ and $v_{1}(t,x,y),v_{2}(t,x,y)$ are the corresponding eigenvectors, then the blue vectors are the eigenvector $v_{1}(t,x,y)$. Then the orange contour in the figure denotes how much light passes through the liquid crystal at a given point in space, which is determined by the average orientation of the liquid crystal molecules $v_{1}(t,x,y)$ and the measure of alignment of molecules in the distribution with the average, $\lambda_{1}(t,x,y)$. Specifically, the color is given by a color mapping with the scalar quantity $\lvert\lambda_{1}(t,x,y)(v_{1}(t,x,y)\cdot(1,0))\rvert,$ so if the molecules are in the vertical direction or are in the isotropic liquid phase (that is, they are not aligned with the average orientation so that $\lambda_{1}(t,x,y)\approx 0$) then light does not pass through the liquid crystal. We see that qualitatively, the high fidelity, DEIM, and NEIM solutions all have similar contour plots, and the blue director fields are also similar. The contour for DEIM appears more similar to the high fidelity contour than NEIM compared to the high fidelity contour for time $t=0.375$, but NEIM is more similar at the final time $t=2$. In both cases, the blue director fields in the DEIM and NEIM plots well-approximate the high fidelity director field, with each plot in the equilibrium state of all horizontal directors.
@@ Acknowledgements
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE 2146752. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. MH also acknowledges the ThinkSwiss Research Scholarship and the EPFL Excellence Research Scholarship. FP acknowledges the “GO for IT” program within the CRUI fund for the project “Reduced order method for nonlinear PDEs enhanced by machine learning”, and the support by Gruppo Nazionale di Calcolo Scientifico (INdAM-GNCS). This work has been conducted within the research activities of the consortium iNEST (Interconnected North-East Innovation Ecosystem), Piano Nazionale di Ripresa e Resilienza (PNRR) - Missione 4 Componente 2, Investimento 1.5 - D.D. 1058 23/06/2022, ECS00000043, supported by the European Union’s NextGenerationEU program.